OpenStudy (anonymous):

2 over 2/5 - 2 over 5/2

OpenStudy (anonymous):

21/5

OpenStudy (anonymous):

So, we have to fractions, what we must do is multiply the numerator of each fraction by the reciprocal of the each fraction's corresponding denominator.

OpenStudy (anonymous):

That is to say: 2*5/2-2*2/5

OpenStudy (anonymous):

which gives us=10/2-4/5 simplfy further=5-4/5

OpenStudy (anonymous):

The lcd becomes 5, and the problem is now=25/5-4/5=21/5